# Kaleidoscopic Scintillation Event Imaging

### v1

@@ Abstract
We propose a kaleidoscopic scintillator to increase light collection in a single-photon camera while preserving the event’s spatial information. The kaleidoscopic geometry creates mirror reflections of the event in known locations for a given event location that are captured by the camera. We introduce theory for imaging an event in a kaleidoscopic scintillator and an algorithm to estimate the event’s 3D position. We find that the kaleidoscopic scintillator design provides sufficient light collection to perform high-resolution event measurements for advanced radiation imaging techniques using a commercial CMOS single-photon camera. Code and data are available at https://github.com/bocchs/kaleidoscopic_scintillator.
@@ 1 Introduction
Figure 1: Method overview. An image is composed of light that is emitted from the scintillation event and reaches the camera either directly or after reflecting off mirrors of a kaleidoscopic scintillator. The spatial relationship between the event and mirror reflections is embedded in a Gaussian mixture model whose likelihood is maximized to estimate the event’s location using the EM algorithm.
High energy particles are created in nuclear processes including from man-made devices and natural interactions. There is a wide variety of particles with varying properties. Detecting and characterizing them is crucial for a broad range of applications including nuclear security [vetter2018gamma], nuclear reactor and stockpile imaging [beaumont2015high], medical imaging [gonzalez2021evolution], archeology [menichelli2007scintillating, ryzewski2013neutron], and astronomy [schonfelder1984imaging]. A particularly useful property of these particles is that many of them can penetrate through dense materials and therefore allow us to image through barriers and inside solid structures. The same penetrating properties, however, make it challenging to build a camera that can perform imaging or vision using high energy particles. A detector needs to have a large volume and density for a particle to interact. At the same time, we would like to know the location and shape of the interaction if it happens. One common approach to achieve this is by using a scintillator.
@@
This work focuses on optimizing imaging geometry in designs that estimate the 3D position of individual scintillation events with a single-photon camera. Currently, such designs estimate an event’s 3D position using depth from defocus and perspective projection [bocchieri2024scintillation]. These designs suffer from low light collection and are susceptible to noise from dark counts. High-speed, single-photon sensors are required for tasks that measure individual events due to high particle incidence rates and the finite number of photons emitted by an event. For these kinds of tasks, increasing exposure time does not increase light collection per event and decreases SNR, which makes maximizing light throughput crucial. Examples where such tasks are performed include the Compton camera [parajuli2022development] and neutron scatter camera [weinfurther2018model], which are devices that measure double-events from individual gamma-rays or neutrons to localize the radiation source.
We propose a kaleidoscopic scintillator geometry with specular surfaces to increase light collection in a camera while preserving the event’s spatial information. A kaleidoscope consists of planar mirrors oriented like a pyramid frustum that contains the object being imaged. Kaleidoscopes and light traps have been studied for imaging and reconstructing extended objects with abundant light [reshetouski2011three, xu2018trapping]. In this paper, we investigate the kaleidoscope as a technique for increasing light collection and reconstructing a point source in a photon-starved environment. We study the kaleidoscope in the context of scintillation event imaging and pose a radiation detection problem as a computer vision problem.
@@ 2 Related Work
Kaleidoscopic vision. Mirrors and kaleidoscopic designs have been used for stereo vision [nene1998stereo, gluckman1999planar, gluckman2002rectified] and multi-view 3D reconstructions [reshetouski2011three, ahn2021kaleidoscopic, ahn2023neural, kawahara2023teleidoscopic, fan2025kaleidoscopic, takahashi2021structure, bangay2004kaleidoscope, mitsumoto19923] of an extended object using one picture from a single camera. A “light-trap” design consists of mirrors oriented such that light entering the trap reaches nearly every position inside the trap [xu2018trapping]. Time-of-flight is used to reconstruct a 3D object inside the trap. In fact, the pyramid-shaped light-trap was found to provide the best object coverage, which is the same shape as the kaleidoscope we use in this work.
Scintillation event measurement designs. Numerous designs exist for measuring scintillation events. Sensors such as silicon photomultiplier (SiPM) arrays or photomultiplier tubes (PMT’s) can be coupled to the surface of a scintillator that is monolithic or pixelated with internal waveguides [gonzalez2021evolution, kawula2021sub, hosokoshi2019development, lee2020development, kataoka2013handy]. Camera designs measure events in a monolithic scintillator by imaging the scintillator with a lens [bocchieri2024scintillation, losko2021new, gustschin2024event, yamamoto2023development, d2021novel, gao2023novel, pleinert1997design, baker2014scintillator, adams2017gamma, balasubramanian2022x]. Using a camera to estimate the 2D position of an event in a thin scintillator, not including depth, is performed using center of mass algorithms [hussey2017neutron, losko2021new, gustschin2024event].
There is limited work that estimates the 3D position of individual events with a camera. One work attempts to remove dark counts and estimate depth from defocus [bocchieri2024scintillation]. Another work uses stereo vision with two objectives and mirrors inside the imaging system to project two views of a cubic scintillator onto a hybrid photon detector for reconstructing the tracks of relativistic electrons [filipenko2014three]. Our proposed system provides multiple views, is more efficient, and uses a commercially available SPAD camera. A plenoptic SPAD camera was proposed to reconstruct events [dieminger2025ultrafast]. The plenoptic camera design’s experimental demonstration was limited by low light collection, obtaining up to 4 photons in an image, while its simulated reconstruction of neutrino tracks achieved high resolution.
@@ 3 Kaleidoscopic event imaging theory
Without loss of generality, we consider a square pyramid scintillator throughout this paper. Each face of the scintillator, except for the base, is a specular surface. The scintillator has an index of refraction $n>1$ and height $h$.
@@ 3.1 Imaging configuration and model
The world coordinate system’s origin is set at the pyramid’s apex with the $z$-axis directed perpendicular toward the pyramid’s base surface. The camera coordinate system’s origin is at the center of the lens with its $z$-axis directed toward the scintillator in the opposite direction of the world’s $z$-axis. The $x$ and $y$ axes among the two coordinate systems are in the same direction. We denote a $z$-coordinate in the world coordinate system as $z_{w}$ and camera coordinate system as $z_{c}$. Transforming between the two coordinate systems is done by subtracting the $z$-coordinate from the lens’ $z$-coordinate in the world coordinate system. We use a thin lens to model the camera. Three important planes to note are the focal plane, the thin lens plane, and the sensor plane. The focal plane is set at the apparent depth of the pyramid’s apex at $z_{w}=h-h/n$. A thin lens with diameter $A$ is placed at a distance $S_{1}$ from the focal plane. The sensor is placed at a distance $S_{2}$ from the lens. The lens’ focal length is set to $f=(S_{1}^{-1}+S_{2}^{-1})^{-1}$. These parameters are illustrated in Fig. 2.
@@
Figure 2: Imaging parameters and coordinate systems. The figure only shows light emitted directly from the event to the camera.
@@ 3.2 Mirror reflections and apertures
is the mirror’s transformation, $\bm{I}$ is the identity matrix, and $\bm{n_{k}}$ is the mirror’s normal vector. The mirror reflection of an event is also a point source of light. The captured image of the mirror reflection is obtained from the photons that reflect off the mirror and into the camera, exhibiting the same defocus blur as if an event were located at $\bm{p_{k}}$. However, due to the finite mirror size, the image on the sensor may be truncated along lines corresponding to the mirror’s edges. This occurs when $\bm{p_{k}}$ is behind another mirror from the camera’s perspective. The photons that are truncated from the image are those that reflect off the mirror adjacent to mirror $k$ near the shared edge, as shown in Fig. 3. Essentially, mirror $k$ behaves like an aperture to a light source at $\bm{p_{k}}$. We denote the area on the sensor where photons from a mirror reflection cannot arrive as the “truncation zone” and its complement as the “acceptance zone”. A line that separates the truncation zone and acceptance zone is a “truncation line”. A 3D visualization of light truncation is shown in Fig. 3. Supplementary Fig. 6 shows a 2D view of the propagation of light and how truncations form in various scenarios. Truncation lines for mirror reflections are derived in Appendix A.
@@
Figure 3: Image truncation example in 3D. The image of an event and one mirror reflection is shown. Light corresponding to the mirror reflection is truncated at the mirror’s edge, resulting in a truncation line in the image. The truncation line only applies to that mirror reflection.
@@ 4.2 Optimization algorithm
5 Experiments
Figure 5: Selected experimental images. Experimental images overlaid with the algorithm’s estimated Gaussian components. Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes. The number of counts in the image and the algorithm’s estimated event location are shown in each image.
@@ 5 Experiments
6 Comparison to prior art
Simulations. Since scintillation events are difficult to produce at controlled locations, direct evaluation of the method requires simulated data. Here, we simulate our system and compare its localization performance to prior 3D localization methods, which estimate depth from defocus in the absence of mirror reflections [bocchieri2024scintillation]. We evaluate position accuracy and resolution over a grid of event locations and over different levels of event brightness. We use the same scintillator geometry, index of refraction, calibrated camera parameter values, and sensor array specifications as in the experiments.
@@ 6 Comparison to prior art
Noise-removing [bocchieri2024scintillation] 0.64 1.04 1.02 1.04
@@
Noise-removing [bocchieri2024scintillation] 0.68 1.05 1.03 1.06
@@
Noise-removing [bocchieri2024scintillation] 0.80 1.20 1.13 1.14
@@
We also test a prior method on non-kaleidoscopic images that attempts to classify and remove dark counts and then directly solve for the event location’s maximum likelihood estimate without weighting photons [bocchieri2024scintillation]. More details of this method are in Appendix G. The algorithms that are applied to non-kaleidoscopic images estimate depth from defocus.
@@ 7 Discussion
Experimental validation of kaleidoscopic image model. Various examples where the Gaussian components predicted by the algorithm coincide with the photon clusters of the event and mirror reflections are shown in Fig. 5. This demonstrates that mirror reflections are indeed being captured in accordance with the proposed model. Examples include events occuring at high (Fig. 5a) and low (Fig. 5b) $z_{w}$-coordinates, mirror reflections lying outside the FOV (Fig. 5c,d), and mirror reflections that are completely truncated (Fig. 5e,f).
@@
Comparison to prior art. The results in Tab. 1 show that the kaleidoscopic design improves 3D localization performance compared to estimating depth from defocus and attains sub-millimeter resolution across all brightness levels tested. The kaleidoscopic design’s performance retains high accuracy and resolution even at lower brightness levels, demonstrating robustness to low photon counts and dark counts.
Limitations and future work. In this work, we only test one specific kaleidoscopic geometry with first-order reflections from one event. Future work may test different kaleidoscopic geometries and higher-order reflections. Developing a method to measure multiple events in one image would allow to construct a Compton camera or neutron scatter camera to localize the radiation source [parajuli2022development, weinfurther2018model].
@@ Appendix G Denoising and estimating depth from defocus (prior method)
Prior methods that estimate a 3D event location using depth from defocus have also attempted to remove dark counts using a denoising algorithm before solving for the location’s maximum likelihood estimate [bocchieri2024scintillation]. The denoising algorithm consists of computing the minimum spanning tree among all photons in an image and removing edges longer than a chosen threshold, $T_{edge}$. The photons in the largest connected component by cardinality are classified as the scintillation photons, and the rest are discarded as dark counts. We test this method on the simulated non-kaleidoscopic images (described in Sec. 6 of the main paper) and report results in Tab. 2. We observe similar performance as the non-kaleidoscopic version of our proposed algorithm.
@@
Table 2: Prior method all simulation results. Estimating depth from defocus on non-kaleidoscopic images after attempting to remove dark counts as in a prior method [bocchieri2024scintillation].
@@ Appendix H Additional figures
Figure 11: Regularization example. An experimental image overlaid with the algorithm’s estimated Gaussian components with (a) $\lambda=0$ (no regularization) and (b) $\lambda=10$ (regularization). Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. The number of counts in the image and the algorithm’s estimated event location are shown in each image. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes.
Figure 12: Selected experimental calibration image. The image is overlaid with the Gaussian components found after manually adjusting the camera parameters. Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes.
Figure 13: Additional selected experimental images. Experimental images overlaid with the algorithm’s estimated Gaussian components. Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes. The number of counts in the image and the algorithm’s estimated event location are shown in each image.
@@
Figure 16: The kaleidoscopic scintillator used in experiments.
Figure 17: Two views of the experimental setup without the gamma-ray source.

### v2

@@ Abstract
We propose a kaleidoscopic scintillator to increase light collection in a single-photon camera while preserving the event’s spatial information. The kaleidoscopic geometry creates mirror reflections of the event in known locations for a given event location that are captured by the camera. We introduce theory for imaging an event in a kaleidoscopic scintillator and an algorithm to estimate the event’s 3D position. We find that the kaleidoscopic scintillator design provides sufficient light collection to perform high-resolution event measurements for advanced radiation imaging techniques using a commercial CMOS single-photon camera.
@@ 1 Introduction
Figure 1: Method overview. An image is composed of light emitted from the scintillation event that reaches the camera either directly or after reflecting off mirrors of a kaleidoscopic scintillator. The spatial relationship between the event and mirror reflections is embedded in a Gaussian mixture model whose likelihood is maximized to estimate the event’s location using the EM algorithm.
High energy particles are created in nuclear processes including from man-made devices and natural interactions. There is a wide variety of particles with varying properties. Detecting and characterizing them is crucial for a broad range of applications including nuclear security [34], nuclear reactor and stockpile imaging [7], medical imaging [16], archeology [25, 31], and astronomy [32]. A particularly useful property of these particles is that many of them can penetrate through dense materials and therefore allow us to image through barriers and inside solid structures. The same penetrating properties, however, make it challenging to build a camera that can perform imaging or vision using high energy particles. A detector needs to have a large volume and density for a particle to interact. At the same time, we would like to know the location and shape of the interaction if it happens. One common approach to achieve this is by using a scintillator.
@@
This work focuses on optimizing imaging geometry in designs that estimate the 3D position of individual scintillation events with a single-photon camera. Currently, such designs estimate an event’s 3D position using depth from defocus and perspective projection [8]. These designs suffer from low light collection and are susceptible to noise from dark counts. High-speed, single-photon sensors are required for tasks that measure individual events due to high particle incidence rates and the finite number of photons emitted by an event. For these kinds of tasks, increasing exposure time does not increase light collection per event and decreases SNR, which makes maximizing light throughput crucial. In general, large, thick scintillators are desirable because they increase the likelihood of events. However, estimating the 3D position of events in a larger volume is more challenging. Applications where 3D event locations are measured include the Compton camera [28], neutron scatter camera [35], positron emission tomography (PET), and single-photon emission computed tomography (SPECT) [16].
We propose a kaleidoscopic scintillator geometry with specular surfaces to increase light collection in a camera while preserving the event’s spatial information. A kaleidoscope consists of planar mirrors oriented like a pyramid frustum that contains the object being imaged. Kaleidoscopes and light traps have been studied for imaging and reconstructing extended objects with abundant light [30, 36]. In this paper, we investigate the kaleidoscope as a technique for increasing light collection and reconstructing a point source in a photon-starved environment. We study the kaleidoscope in the context of scintillation event imaging and pose a radiation detection problem as a computer vision problem.
@@ 2 Related Work
Kaleidoscopic vision. Mirrors and kaleidoscopic designs have been used for stereo vision [27, 14, 15] and multi-view 3D reconstructions [30, 3, 2, 21, 11, 33, 6, 26] of an extended object using one picture from a single camera. A “light-trap” design consists of mirrors oriented such that light entering the trap reaches nearly every position inside the trap [36]. Time-of-flight is used to reconstruct a 3D object inside the trap. In fact, the pyramid-shaped light-trap was found to provide the best object coverage, which is the same shape as the kaleidoscope we use in this work.
Scintillation event measurement designs. Numerous designs exist for measuring scintillation events. Sensors such as silicon photomultiplier (SiPM) arrays or photomultiplier tubes (PMT’s) can be coupled to the surface of a scintillator that is monolithic or pixelated with internal waveguides [16, 22, 18, 23, 20]. Camera designs measure events in a monolithic scintillator by imaging the scintillator with a lens [8, 24, 17, 37, 9, 13, 29, 4, 1, 5]. Using a camera to estimate the 2D position of an event in a thin scintillator, not including depth, is performed using center of mass algorithms [19, 24, 17].
There is limited work that estimates the 3D position of individual events with a camera. One work attempts to remove dark counts and estimate depth from defocus [8]. Another work uses stereo vision with two objectives and mirrors inside the imaging system to project two views of a cubic scintillator onto a hybrid photon detector [12]. Our proposed system provides multiple views, is more efficient, and uses a commercially available CMOS single-photon camera. A plenoptic SPAD camera was proposed to reconstruct events [10]. The plenoptic camera design’s experimental demonstration was limited by low light collection, obtaining up to 4 photons in an image, while its simulated reconstruction of neutrino tracks achieved high resolution.
@@ 3 Kaleidoscopic Event Imaging Theory
We consider a square pyramid scintillator throughout this paper. Each face of the scintillator, except for the base, is a specular surface. The scintillator has an index of refraction $n>1$ and height $h$.
@@ 3.1 Imaging Configuration and Model
The world coordinate system’s origin is set at the pyramid’s apex with the $z$-axis directed perpendicular toward the pyramid’s base surface. The camera coordinate system’s origin is at the center of the lens with its $z$-axis directed toward the scintillator in the opposite direction of the world’s $z$-axis. The $x$ and $y$ axes among the two coordinate systems are in the same direction. We denote a $z$-coordinate in the world coordinate system as $z_{w}$ and camera coordinate system as $z_{c}$. Transforming between the two coordinate systems is done by subtracting the $z$-coordinate from the lens’ $z$-coordinate in the world coordinate system. We use a thin lens to model the camera. Three planes to note are the focal plane, thin lens plane, and sensor plane. The focal plane is set at the apparent depth of the pyramid’s apex at $z_{w}=h-h/n$. A thin lens with diameter $A$ is placed at a distance $S_{1}$ from the focal plane. The sensor is placed at a distance $S_{2}$ from the lens. The lens’ focal length is set to $f=(S_{1}^{-1}+S_{2}^{-1})^{-1}$. Parameters are illustrated in Fig. 2a.
@@
Figure 2: (a) Imaging parameters and coordinate systems. Only light emitted directly from the event to the camera is shown. (b) Image truncation example in 3D. The image of an event and one mirror reflection is shown. Light corresponding to the mirror reflection is truncated at the mirror’s edge, resulting in a truncation line in the image. The truncation line only applies to that mirror reflection.
@@ 3.2 Mirror Reflections and Apertures
is the mirror’s transformation, $\bm{I}$ is the identity matrix, and $\bm{n_{k}}$ is the mirror’s normal vector. The mirror reflection of an event is also a point source of light. The captured image of the mirror reflection is obtained from the photons that reflect off the mirror and into the camera, exhibiting the same defocus blur as if an event were located at $\bm{p_{k}}$. However, due to the finite mirror size, the image on the sensor may be truncated along lines corresponding to the mirror’s edges. This occurs when $\bm{p_{k}}$ is behind another mirror from the camera’s perspective. The photons that are truncated from the image are those that reflect off the mirror adjacent to mirror $k$ near the shared edge, as shown in Fig. 2b. Essentially, mirror $k$ behaves like an aperture to a light source at $\bm{p_{k}}$. We denote the area on the sensor where photons from a mirror reflection cannot arrive as the “truncation zone” and its complement as the “acceptance zone”. A line that separates the truncation zone and acceptance zone is a “truncation line”. A 3D visualization of light truncation is shown in Fig. 2b. Supplementary Fig. 6 shows a 2D view of the propagation of light and how truncations form in various scenarios. Truncation lines are derived in Appendix A.
@@ 4.2 Optimization Algorithm
5 Hardware Experiments
Figure 4: Experimental setup. (a) The setup without the gamma-ray source. (b) The kaleidoscopic scintillator.
Figure 5: Selected experimental images. Experimental images overlaid with the algorithm’s estimated Gaussian components. Each dashed circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Red and blue circles represent the event and mirror reflections, respectively. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes. The number of counts in the image and the algorithm’s estimated event location are shown in each image.
@@ 5 Hardware Experiments
6 Simulations
Comparison to prior art. Since scintillation events are difficult to produce at controlled locations, direct evaluation of the method requires simulated data. Here, we simulate our system and compare its localization performance to prior 3D localization methods, which estimate depth from defocus in the absence of mirror reflections [8]. We evaluate position accuracy and resolution over a grid of event locations and over different levels of event brightness. We use the same scintillator geometry, index of refraction, calibrated camera parameter values, and sensor array specifications as in the experiments.
@@ 6 Simulations
Noise-removing [8] 0.64 1.04 1.02 1.04
@@
Noise-removing [8] 0.68 1.05 1.03 1.06
@@
Noise-removing [8] 0.80 1.20 1.13 1.14
@@
We also test a prior method on non-kaleidoscopic images that attempts to classify and remove dark counts and then directly solve for the event location’s maximum likelihood estimate without weighting photons [8]. More details of this method are in Appendix G. The algorithms that are applied to non-kaleidoscopic images estimate depth from defocus.
@@ 7 Discussion
Experimental validation of kaleidoscopic image model. Various examples where the Gaussian components predicted by the algorithm coincide with the photon clusters of the event and mirror reflections are shown in Fig. 5. This demonstrates that mirror reflections are indeed being captured in accordance with the proposed model. Examples include events occurring at high (Fig. 5a) and low (Fig. 5b) $z_{w}$-coordinates, mirror reflections lying outside the FOV (Fig. 5c,d), and mirror reflections that are completely truncated (Fig. 5e,f).
@@
Simulations. The results in Tab. 1 show that the kaleidoscopic design improves 3D localization performance compared to estimating depth from defocus and attains sub-millimeter resolution across all brightness levels tested. The kaleidoscopic design’s performance retains high accuracy and resolution even at lower brightness levels, demonstrating robustness to low photon counts and dark counts.
Limitations and future work. In this work, we only test one specific kaleidoscopic geometry with first-order reflections from one event. Future work may test different kaleidoscopic geometries and higher-order reflections. Developing a method to measure multiple events in one image would allow to construct a Compton camera or neutron scatter camera to localize the radiation source [28, 35].
@@ Acknowledgments
This material is based upon work supported by the Department of Energy / National Nuclear Security Administration under Award Number(s) DE-NA0004196.
@@ Appendix G Denoising and Estimating Depth from Defocus (Prior Method)
Prior methods that estimate a 3D event location using depth from defocus have also attempted to remove dark counts using a denoising algorithm before solving for the location’s maximum likelihood estimate [8]. The denoising algorithm consists of computing the minimum spanning tree among all photons in an image and removing edges longer than a chosen threshold, $T_{edge}$. The photons in the largest connected component by cardinality are classified as the scintillation photons, and the rest are discarded as dark counts. We test this method on the simulated non-kaleidoscopic images (described in Sec. 6 of the main paper) and report results in Tab. 2. We observe similar performance as the non-kaleidoscopic version of our proposed algorithm.
@@
Table 2: Prior method all simulation results. Estimating depth from defocus on non-kaleidoscopic images after attempting to remove dark counts as in a prior method [8].
@@ Appendix H Additional Figures
Figure 11: Regularization example. An experimental image overlaid with the algorithm’s estimated Gaussian components with (a) $\lambda=0$ (no regularization) and (b) $\lambda=10$ (regularization). Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Red and blue circles represent the event and mirror reflections, respectively. The number of counts in the image and the algorithm’s estimated event location are shown in each image. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes.
Figure 12: Selected experimental calibration image. The image is overlaid with the Gaussian components found after manually adjusting the camera parameters. Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Red and blue circles represent the event and mirror reflections, respectively. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes.
Figure 13: Additional selected experimental images. Experimental images overlaid with the algorithm’s estimated Gaussian components. Each dashed red circle is centered on the Gaussian component’s mean. The inner and outer circles are one and two standard deviations in radius, respectively. Red and blue circles represent the event and mirror reflections, respectively. Pixels with a photon are enlarged with a $3\times 3$ filter for visualization purposes. The number of counts in the image and the algorithm’s estimated event location are shown in each image.